2000

Progress report: the site in Spring/Summer 2000

January
The first double triple-pendulum suspension (i.e. triple-pendulum suspension for mirror and for reaction mass) has been installed for the near mirror in TCE. The suspension in the east end house has been successfully tested for vacuum suitability.

February
We are starting the installation of the suspension in the power recycling tank. The suspension in TCE has been successfully tested for vacuum suitability. We performed a first test of the data acquisition system by taking data for 8 h from the locked mode cleaner.

March
A horizontal motor has been installed for the length control of mode cleaner 2. A change in control topology has been introduced in order to improve control of the mode-cleaner lengths and of the laser frequency: Instead of a reference cavity the mode cleaners serve as a frequency reference for the laser pre-stabilization. The optimum transmission of the modulation side bands of the power-recycling lock of mode cleaner 2 serves as a dc fequency reference.

April
The new control concept has been tested successfully. The shadow-sensor/coil units of the local controls have been redisigned.

May
The first main mirror (test optics) has been installed in the east end tank TFE. It has been pre-adjusted by means of a He-Ne laser being collinearly reflected through the 600 m arm length.
The active seismic isolation has been demonstrated for the first time in TCE. The horizontal and vertical components of the seismic excitation have been suppressed at the same time by about a factor of ten in the fourier-frequency range of the pendulum resonances.

June
The pockels cells have been replaced by new ones made of LiNbO having a higher modulation index.
During the first half of this year the LabView control programs have been installed. The analog local control electronics has been replaced step by step by LabView controlled ones.

 
Go to Editor View
loading content